# Tradeoffs Between Alignment and Helpfulness in Language Models

### v2

@@ 5 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how it changes as we increase the vector norms of representation engineering. The experiments were performed with Llama 2 13B chat and Llama 2 13B Touvron et al. (2023). For the aligned Llama 2 13B chat, we calculate the behavior expectation w.r.t.behaviors "harmfulness" and "racism", as a function of representation engineering vector coefficients using vectors for the corresponding behaviors. For the misaligned Llama 2 13B, we do the same w.r.t."harmlessness". Then, we calculate the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Additional experimental details can be found in appendix D.
@@
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t.a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the "harmfulness" behavior on Llama 2 13B chat, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For "racism" we used biased and unbiased statements from the StereoSet dataset Nadeem et al. (2020). The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
@@
To calculate the racism behavior expectation, we used a behavior scoring function that assigns an answer $B(answer)=\pm 1$ to agreeing/disagreeing with a racist statement, and calculated the expectation value of this function w.r.t.the model distribution, which amounts to the difference between probabilities of agreeing and disagreeing with a racist statement. To calculate harmfulness behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model agrees/refuses to answer a harmful instruction and calculated its expectation value, which amounts to the difference between probabilities of agreeing and disagreeing to answer.
Figure 4(a) shows racism behavior expectation as a function of bias PCA vector coefficients injected into the model. Figure 4(b) shows harmfulness behavior expectation as a function of harmful PCA vector coefficients injected into the model. Overall we see that on both behaviors, the behavior expectation changes like a hyperbolic tangent, as expected of theorem 1.
@@
Figure 4: (a) Racism behavior expectation of Llama 2 13B chat as a function of coefficient of injected bias PCA vectors. (b) Harmful behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors.
Note that the behavior scoring functions defined here are positive for behaviors that are negative. This was done because the model used is already aligned, so we demonstrate a flip in its behavior from positive to negative, while in Zou et al. (2023a) they show a flip in the opposite direction. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, it suffices to flip the behavior in one direction.
Still, we also perform the reversed experiment, of aligning a pretrained model, Llama 2 13B w.r.t."harmlessness" with representation engineering. To this end, we extracted harmful instructions from AdvBench Robey et al. (2021; 2022) and created pairs of instruction+harmful/unharmful responses. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
To calculate harmfulness behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model refuses/agrees to answer a harmful instruction and calculated its expectation value, which amounts to the difference between the difference between probabilities of disagreeing and agreeing to answer.
Figure 5 shows harmlessness behavior expectation as a function of bias PCA vector coefficients injected into the model. As can be seen, the behavior expectation changes as a hyperbolic tangent, as expected of theorem 1.
Figure 5: Harmlessness behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors.
To calculate helpfulness, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) and calculated the probability that the model assigns the correct one. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 6a shows this for the "fairness" behavior vectors and 6b shows this for the "harmfulness" behavior vectors. As can be seen, the probability to answer correctly starts by first dropping slowly then decays to zero, similarly to the proposed bound in theorem 2. The average helpfulness is usually parabolic around $r_{e}=0$ but sometimes has a plateau instead, which corresponds to the helpfulness for different queries not being peaked exactly at $r_{e}=0$, for them the bound is not tight, hence a decrease that is not parabolic around $r_{e}=0$ can be observed. Still, the slow decrease of helpfulness for small coefficients of representation engineering is seen in those cases as well.
In figure 7, we perform the same experiment when restricting the probabilities to the answers of the multiple choice question, A,B,C,D. The same result of theorem 2 is observed. Moreover, we see that the average helpfulness at the tails saturates at $\approx 1/4$, according to corollary 2.
Figure 6: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Figure 7: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b). The probability of correct answer was measured relative to the answers A, B, C, D.
Similarly, we calculate the helpfulness for Llama 2 13B as a function of coefficient of the PCA vectors used to align the model for harmlessness. Figure 8(a) shows that the probability to answer correctly starts by dropping slowly then decays to zero, similarly to the proposed bound in theorem 2. Figure 8(b) shows this when restricting the answers to "A", "B", "C", "D".
Figure 8: Probability of correct answer of Llama-2-13B on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted to align the model for harmlessness. (a) On entire vocabulary (b) Answers restricted to "A", "B", "C", "D".
@@ 6 Discussion
We show the dynamics of representation editing. On the one hand a linear increase in the desired behavior and on the other hand a parabolic decrease on general performance of basic tasks.
There are two limitations to the terms "linear alignment" and "quadratic helpfulness". Notice that since behavior expectation is bounded between $-1$ to $+1$, the change happens within a bounded range of coefficients $r_{e}$. It may include $r_{e}=0$ (figure 4a) but it may also be outside its range 4b), such that initially one might have to work harder before seeing the payoff. As explained in section 4, the linear increase will be seen already at $r_{e}=0$ if the model’s behavior is negative but has a small probability for positive behavior as well. Also note that the quadratic bound for helpfulness is an upper bound and as explained may not be tight, indeed the tails observed in figure 6 are more typical of exponential decay than parabolic decay. Still, the interesting observation of a helpfulness decay around $r_{e}=0$ which is initially slower than alignment increase is observed.
Notice that our framework uses the model’s next token prediction to determine alignment and helpfulness. Initially this may seem to restrict the generality of our results, but note that for alignment, it is often the first token that determines whether an answer is aligned or not ("Sorry, as an AI I cannot…" vs "Sure! Here’s…"). For helpfulness this point is more subtle, as some tasks are more intricate, however, as demonstrated by our experiments, there are numerous tasks that can be reduced to next token prediction.
@@
One may argue that these results are applicable only when one behavior vertical is enhanced with representation engineering, and the rest remain untouched leading to a reduction in performance on them. Indeed, one may apply representation engineering in several behavior verticals to improve on many behaviors, however so long as an important vertical is left out, our results imply that the LLM’s performance on it will reduce. Furthermore, one can attempt to apply this framework for analyzing the effects of normal finetuneing on alignment and helpfulness, as it too could be summed up to a change in the representations of the LLM to maximize the likelihood of desired outputs. However, the changes of a finetuning process to a model’s representations are different than representation engineering’s change: While finetuning creates small changes to the model’s representation at each training step on several different behaviors, that sums up to a large overall change, representation engineering takes a large step in one behavioral direction, disregarding other behavior verticals. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 3), unlike a finetuning process where this does not necessarily happen. However, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
Lastly, a limitation of our results for alignment guarantee is that they may rely on injecting vectors of large norms to the model. However, for very large norms, the model becomes unhelpful to the point of outputting meaningless words.
@@
Overall, we hope that our theoretical work will shed light on the mechanism of representation engineering, which provides a new interesting direction for language model alignment.
@@ A.1 Introduction of assumptions
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix E.
@@ A.2 Discussion of assumptions
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[\langle v,r_{good}-r_{bad}\rangle]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix E we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
@@ A.3 Experiments for assumptions
Here we empirically check the validity of our assumptions. The experiments were performed on Llama 2 13B chat.
@@
We show that $\delta r_{e}(q)$ increases linearly for coefficients that are not too large (figure 9). Here we use the above mentioned fairness PCA vectors. Here we average on different datasets from MMLU:
Figure 9: Linear increase in the norm of $U\delta r_{e}(q)$ for small coefficients.
@@
We checked that the change in logit distribution is normal (figure 10) as proposed in assumption 3 and verify that indeed it is approximately normal with the standard deviation growing linearly with the injected vector norm. Note that the latter condition is because in assumption 3, we normalize $\delta r_{e}(q)$ before taking the inner product in equation 12, while here we do not, thus we expect to see the standard deviation of the Gaussians grow linearly with $|\delta r_{e}|$ which as shown above grows linearly with $r_{e}$. In figure 10(a) and (b) we show the density of change in the top ten token’s logits due to representation engineering on the dataset "international law" from MMLU for the PCA vectors of fairness and harmfulness with different coefficients. In figure 10(c) and (d) we show that the standard deviation of the distributions obtained in (a) and (b) increases linearly with the size of the coefficient. Notice that here we show $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$ distributes normally instead of $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, since if the latter distributes normally, then the former is a subtraction of two normal distributions and should also be normal. The benefit of using the former instead of the latter is that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail, which can lead to an increase in the correct answer’s probability as $|\delta r_{e}(q)|$ is increased.
Figure 10: Distribution of the change in token logits minus the logit of the correct answer. As can be seen, it is approximately normal, and the standard deviation grows linearly with the coefficient size, which is linear in $|\delta r_{e}(q)|$.
@@
In figure 11 we plot the center of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min.
Figure 11: Separation between representation clusters of positive and negative behavior tokens for two harmful instructions from the AdvBench dataset. Note that the model is aligned at $r_{e}=0$, thus the common tokens of $r_{e}=0$ are of aligned answers.

### v3

@@ 4 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how they change as we increase the vector norms of representation engineering. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, we use representation engineering to show an increase in alignment of an unaligned pretrained model (specifically we use Llama 2 13B Touvron et al. (2023)), and a decrease in alignment of an aligned RLHF model (we use Llama 2 13B chat Touvron et al. (2023)). For the unaligned model, we calculate the behavior expectation w.r.t. behaviors "harmless" and "not-racist", as a function of representation engineering vector coefficients and show an improvement in alignment. For the aligned model, we do the same for the behaviors "harmful" and "racist" and show a decrease in alignment. Our experiments show an effect of representation engineering on alignment that matches theorem 1. Then, we calculate helpfulness as the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Our experiments demonstrate that helpfulness changes as in theorem 2. Additional experimental details can be found in appendix E.
@@
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t. a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the "harmful" behavior on the aligned model, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For "harmless" behavior on the the unaligned model, the above approach of contrasting positive and negative requests does not work, since the model did not undergo alignment, so it equally agrees to answer both types of requests. Thus, the produced engineered representations never steer the model towards not answering any request. So, inspired by the method of preference learning, we contrast aligned and misaligned responses to harmful instructions from AdvBench. For "racism" on the aligned model, we used biased and unbiased statements from the StereoSet dataset Nadeem et al. (2020). For "not-racist" on the unaligned model, we used the racist statements from above, followed by aligned and misaligned responses, and contrasted them. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
Alignment Measurement:
@@
To calculate harmful behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model agrees/refuses to answer a harmful instruction and calculated its expectation value, which amounts to the difference between probabilities of agreeing and disagreeing to answer. To calculate the racism behavior expectation, we used a behavior scoring function that assigns an answer $B(answer)=\pm 1$ to agreeing/disagreeing with a racist statement, and calculated the expectation value of this function w.r.t.  the model distribution, which amounts to the difference between probabilities of agreeing and disagreeing with a racist statement.
@@
Figure 4: Plots of behavior expectation as a function of the coefficients of representation engineering vectors injected to the model. The blue line is the direct measurement, the orange line is a plot of the bound from theorem 1. (a) Harmless behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (b) Racism behavior expectation of Llama 2 13B as a function of coefficient of injected bias PCA vectors.(c) Harmful behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (d) Racism behavior expectation of Llama 2 13B chat as a function of coefficient of injected bias PCA vectors.
Figure 4a(c) shows harmless (harmful) behavior expectation as a function of harmless (harmful) PCA vector coefficients injected into the misaligned (aligned) model. Figure 4b(d) shows not-racist (racist) behavior expectation as a function of bias PCA vector coefficients injected into the misaligned (aligned) model. Overall we see that on both behaviors and both models, the behavior expectation changes like a hyperbolic tangent, as expected of theorem 1, which can be seen by the fitted curve of the data to a bound of the form of theorem 1 when using $\Delta\lambda$ as a free parameter that fits the measurements. The value of $\Delta\lambda$ corresponding to the curve is $0.5-3$ while our empirically estimated value of $\Delta\lambda$ from the data based on the linear classification condition of the last hidden layer change is $0.1-0.4$ (for details on the empirical estimation see appendix A.3). The difference between these two ranges may be attributed to the method of the empirical estimation of $\Delta$ that looks for an upper bound on it on the entire $r_{e}$ range, while the main change in alignment in figure 4 occurs in a more specific range, where the upper bound of $\Delta$ is evidently bigger. We note that for all behaviors, $r_{e}=2.5$ suffices for a significant change in behavior expectation, taking it from negative to positive. It is left to observe the decrease in helpfulness and verify that it is not too small.
Helpfulness Measurement:
To calculate helpfulness, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) over a variety of domains (international law, medical genetics, high school computer science) and calculated the probability that the model assigns the correct one. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 5a(c) shows this for the harmless (harmful) behavior vectors on the misaligned (aligned) model and 5b(c) shows this for the not-racist (racist) behavior vectors for the misaligned (aligned) model. Here we restricted the probabilities to the answers of the multiple choice question, A,B,C,D.
According to theorem 2 and corollary 2, around $r_{e}=0$, helpfulness should decrease parabolically, which can be seen by Taylor expanding the bounds, yielding $\mathbb{P}_{correct}\approx f(0)+\frac{1}{2}f^{\prime\prime}(0)r_{e}^{2}$ where $f^{\prime\prime}(0)<0$. Then, according to corollary 3, since there are $N=4$ answers to choose from, the probability of the correct answer should on average converge to $1/4$. To demonstrate this behavior in the empirical measurements, we plot a bound of the form of theorem 2 with the boundary conditions of corollary 3 (for further explanation on the theoretical justification of using this bound see appendix E.4). We do so with free parameter $\lambda\sigma\beta$, and $\alpha$ set to $1$ (in our theoretical bound it is smaller, but it is due to the bound being centered at $r_{e}=0$ while the peak is not guaranteed to be, hence it may overshoot), from which we find $\lambda\sigma\beta$ in the range of $0.33-0.66$. With our empirically estimated values of $\lambda\sigma\beta$ from direct measurement of the noise injected due to representation engineering, in the range $0.1-0.4$ (see details on the empirical estimation in appendix A.3).
Figure 5: Helpfulness measurement: the probability assigned to the correct answer to questions from different MMLU tests (international law, medical genetics, high school computer science), as a function of representation engineering vector coefficients injected to the model. Here the probability of the correct answer was measured relative to the answers A, B, C, D. The red line plots the bound of theorem 2 for free parameters on "international law". (a) Harmless behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (b) Racism behavior expectation of Llama 2 13B as a function of coefficient of injected bias PCA vectors. (c) Harmful behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (d) Racism behavior expectation of Llama 2 13B chat as a function of coefficient of injected bias PCA vectors.
Notably, for $r_{e}=2.5$, the decrease in helpfulness is still not too great, while as mentioned previously, alignment is significantly increased. Further note that the decrease in helpfulness is not attributed to the model’s refusal to answer questions, as one might suspect for an injection of harmless vectors. This is because for both positive and negative coefficients the helpfulness drops, while the refusal to answer harmful queries grows only in one direction.
In appendix E, we perform the same experiment but without restricting the answers to A,B,C,D (over the full vocabulary). The same result of theorem 2 is observed, though the helpfulness decays to zero instead of $1/4$ due to $1/|V|$ scaling of corollary 3. The average helpfulness is usually parabolic around $r_{e}=0$ but sometimes has a plateau instead, which corresponds to the helpfulness for different queries not being peaked exactly at $r_{e}=0$, for them the bound is not tight, hence a decrease that is not parabolic around $r_{e}=0$ can be observed.
@@ 5 Discussion
In this work, we study the benefits of representation editing for LLM alignment from a theoretical perspective. We find that increasing the magnitude of the vectors injected to the model leads to improved alignment; we theoretically quantify this improvement as linear in the vectors’ magnitude, and validate our result empirically. A practical outcome of our result is a guarantee of alignment when using the representation engineering method. Such theoretical guarantees cannot be made without altering the model at inference time – Wolf et al. (2023) show that prompt based alignment methods can always be undone. Our result thus crystallizes an inherent advantage of representation engineering over competing alignment methods.
On the other hand, our framework indicates a degradation of the model’s general capabilities when representation engineering is applied. We theoretically quantify this degradation to be parabolic in the injected vectors’ magnitude, which puts a bound on the strength with which representation engineering should be performed to keep the model reliable for different uses. While our theoretical bound is an upper bound on the helpfulness, we observe this parabolic behavior empirically as well.
@@
In principle, our framework may be generalized for theoretically analyzing the effects of normal finetuneing on alignment and helpfulness, as it too could be summed up to a change in the representations of the LLM to maximize the likelihood of desired outputs. In particular, each step in preference learning is equivalent to a representation engineering injection with coefficient that equals to the learning rate (see appendix G). However, we leave this for future work as the changes of a finetuning process to a model’s representations are different than representation engineering’s change: While finetuning creates small changes to the model’s representation at each training step on several different behaviors, that sums up to a large overall change, representation engineering takes a large step in one behavioral direction, disregarding other behavior verticals. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 3), unlike a finetuning process where this does not necessarily happen. Hence in regards of maintaining helpfulness, finetuning has an advantage, however, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
@@
Overall, we hope that our theoretical work will shed light on the mechanism of representation engineering, which constitutes a new interesting direction for language model alignment.
@@ A.1 Introduction of assumptions
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix F.
@@ A.2 Discussion of assumptions
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[|\langle v,r_{good}-r_{bad}\rangle|^{2}]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix F we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
@@ A.3 Experiments for assumptions
Here we empirically check the validity of our assumptions and empirically estimate the values of the parameters in the bounds. The experiments were performed on Llama 2 13B and Llama 2 13B chat. We first verify a linear relation between the representation engineering coefficient $r_{e}$ to the last hidden layer change of assumption 1, which yields $\lambda$. Then, we verify the normal distribution assumption 3 and the linear classification of assumption 2.
@@
For a query $q$ we define $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ as the change of the representation of the query in the final layer. where $r^{(L)}(q,0)$ is the representation if we injected no vector (the default model representation) and $r^{(L)}(q,r_{e})$ is the representation given that we inject a vector of norm $r_{e}$ at a range of layers. We show that the norm of $\delta r_{e}(q)$ increases linearly with $r_{e}$ when $r_{e}$ is not too large (figure 6). Here we use the above mentioned fairness PCA vectors. We average on different queries from a few datasets taken from MMLU.
In practice we look at $U\delta r_{e}(q)$, where $U$ is the transformation taking from the final layer representation to the logits vector. Since this is a linear transformation, showing a linear relationship between $r_{e}$ and $|U\delta r_{e}(q)|$ implies a linear relationship between $r_{e}$ and $|\delta r_{e}(q)|$.
Figure 6: Linear increase in the norm of $U\delta r_{e}(q)$ for small coefficients, when injected with "racist" vectors.
In figures 7 and 8 we plot the change in norm for Llama 2 13B chat (injected with racist vectors) and Llama 2 13B (injected with not racist vectors) respectively, on the datasets "international law", "medical genetics" and "high school computer science". We add fitted curves to estimate $\lambda$. We find that it is in the range $40-60$.
Figure 7: Norm of the final hidden layer representation change as a function of representation engineering coefficient, for Llama 2 13B chat, on different MMLU datasets. The fitted linear curves estimate $\lambda$.
Figure 8: Norm of the final hidden layer representation change as a function of representation engineering coefficient, for Llama 2 13B, on different MMLU datasets. The fitted linear curves estimate $\lambda$.
@@
As proposed in assumption 3, we show here that the projection of a given answer on the representation change $\delta r_{e}(q)$ is approximately normal, with mean 0. (Assuming the question asked is not connected to the property we are changing with the representation engineering). In assumption 3 we looked at the normalized change: $\langle\frac{\delta r_{e}(q)}{\lvert\lvert\delta r_{e}(q)\rvert\rvert},U^{T}e_{i}\rangle$. Here we will look at $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, so we expect the distribution to be:
$\langle\delta r_{e}(q),U^{T}e_{i}\rangle\sim N(0,\sigma^{2}\lvert\lvert\delta r_{e}(q)\rvert\rvert^{2})$
Meaning the standard deviation scales linearly with the norm of $\delta r_{e}(q)$. Since $r_{e}$ scales linearly with $\delta r_{e}(q)$, we expect the standard deviation to also scale linearly with $r_{e}$. In practice we look at $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$, the advantage of this being that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail.
To create the plot, for each question in a dataset, we look at the top 10 answers $e_{i},i\in[10]$ (with no representation engineering). Now for a given $r_{e}$ coefficient, we calculate the projection of these answers on $\delta r_{e}(q)$. We then aggregate these projections for all the questions in a few dataset and look at their histogram and at their standard deviation. We repeat this for different $r_{e}$ norms.
Figure 9: (a) ((b)) Distribution of the change in token logits minus the logit of the correct answer of Llama 2 13B chat when injected with racist (harmful) vectors. As can be seen, it is approximately normal, and in (c) and (d) the standard deviation grows linearly with the coefficient size $r_{e}$, which is linear in $|\delta r_{e}(q)|$.
Figure 10: (a) ((b)) Distribution of the change in token logits minus the logit of the correct answer of Llama 2 13B chat when injected with harmless (not-racist) vectors. As can be seen, it is approximately normal, and in (c) and (d) the standard deviation grows linearly with the coefficient size $r_{e}$, which is linear in $|\delta r_{e}(q)|$.
The tangent of the curve of figure 9c,d is $\lambda\sigma$, as the curve is the standard deviation of $\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot|\delta r_{e}(q)|=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot\lambda r_{e}$, from assumption 1, and the inner product is a normal variable of standard deviation $\sigma$, hence the tangent is $\lambda\sigma$. From the linear curve, we estimate $\lambda\sigma=0.5$, thus $\lambda\sigma\beta\approx 0.8\cdot 0.5$, as it is the mean of a half-normal distribution with parameter $\lambda\sigma$, which is approximately $0.8\lambda\sigma$.
Similarly, for the pretrained model, we find that $\lambda\sigma=0.2$ and $0.1$ for fairness and harmlessness respectively.
@@
And by assumption 1, it is equivalent to:
$\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle\geq\Delta\lambda\cdot r_{e}$ (12)
In figure 11 and 12, we plot the distance between the centers of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased, for Llama 2 13B chat and Llama 2 13B respectively. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve of the form in equation 12, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min. Thus by equation 12, the tangent of the lower bounding lines of figures 11 and 12 are an estimate for $\Delta\lambda$. From, this we get that $\Delta\lambda$ is approximately $0.1-0.3$
Figure 11: Separation between representation clusters of positive and negative behavior tokens induced by $\delta r_{e}(q)$ on Llama 2 13B chat for three harmful instructions from the AdvBench dataset.
Figure 12: Separation between representation clusters of positive and negative behavior tokens induced by $\delta r_{e}(q)$ on Llama 2 13B for three harmful instructions from the AdvBench dataset.
In practice, the good and bad tokens were chosen beforehand as the top 40 tokens of the models when representation engineering is applied and when it is not applied (meaning in one case the model is aligned and in the other it is not).